8^n/3 = 1/4 then what is the value of n^n ?
a -4 b -1/2 c - 1/4 d 1/4 e 1/2

Answer by stanbon(75887)

8^n/3 = 1/4 then what is the value of n^n ?
---------------------
Raise both sides to the 3rd power to get:
8^n = 1/64
---
8^n = 8^-2
---
n = -2
---
Then (n^n) = (-2)^(-2) = 4^-1 = 1/4
=========================================
Cheers,
Stan H.
============

a -4 b -1/2 c - 1/4 d 1/4 e 1/2